# Grokking in LLM Pretraining? Monitor Memorization-to-Generalization without Test

### v3

@@ Abstract
This paper presents the first study of grokking in practical LLM pretraining. Specifically, we investigate when an LLM memorizes the training data, when its generalization on downstream tasks starts to improve, and what happens if there is a lag between the two. Unlike existing works studying when a small model generalizes to limited and specified tasks during thousands epochs’ training on algorithmic data, we focus on a practical setting for LLMs, i.e., one-epoch pretraining of next-token prediction on a cross-domain, large-scale corpus, and generalization on diverse benchmark tasks covering math/commonsense reasoning, code generation, and domain-specific retrieval. Our study, for the first time, verifies that grokking still emerges in pretraining mixture-of-experts (MoE) LLMs, though different local data groups may enter their grokking stages asynchronously due to the heterogeneity of their distributions and attributions to others. To find a mechanistic interpretation of this local grokking, we investigate the dynamics of training data’s pathways (i.e., expert choices across layers in MoE). Our primary discovery is that the pathways evolve from random, non-smooth across layers, instance-specific to more structured and transferable across samples, despite the converged pretraining loss. This depicts a transition from memorization to generalization. Two novel metrics are developed to quantify these patterns: one computes the pathway similarity between samples, while the other measures the consistency of aggregated experts between subsequent layers for each sample. These training data based metrics induce zero cost but can faithfully track and monitor the generalization of LLMs on downstream tasks, which, in conventional settings, requires costly instruction tuning and benchmark evaluation. We also ground our findings in a theoretical analysis of one-layer MoE, showing that more structured pathways improve the generalization bound.
@@ 1 Introduction
  • Grokking still occurs during the one-epoch pretraining of practical-scale LLMs, but it is local and asynchronous for different data groups, unlike global grokking for all data in previous works.
@@ 3 Grokking (Delayed Generalization) in LLM Pretraining
  • One-epoch pretraining: Most LLMs’ pretraining only takes one epoch, so the model has been trained only once on each sample, and its training loss may even converge before visiting later data. However, such “memorization” of unseen data is not caused by overfitting or repeated replay. Instead, it reflects the interpolation or composition of learned data, which might be counted as “generalization” in previous works but differs from the emergent generalization of LLMs on new tasks.
@@ 3.1 Study Training Dynamics of LLM Pretraining
Since $\ell_{i}(\theta)$ directly measures how the model memorizes each token in $x_{i}$, we will track it in the pretraining process and identify the sample with consistently small and converged loss across checkpoints as memorized. Specifically, for a sample $x_{i}$ with loss $\ell_{i}(\theta_{t})$ at pretraining step $t$, we identify it as memorized since step $t_{i}^{*}$ if $\ell_{i}(\theta_{t})\leq\varepsilon$ and $|\ell_{i}(\theta_{t})-\ell_{i}(\theta_{T})|\leq\delta$ for all $t\geq t_{i}^{*}$, where $\varepsilon$ is a small threshold, $\delta$ enforces stability over time, and $T$ indexes the final pretraining checkpoint.
@@ 3.2 Asynchronous Memorization, Delayed Generalization, & Local Grokking
These observations reveal the complex training dynamics of LLMs caused by training data heterogeneity and their uneven attributions to different benchmark tasks. It implies that the widely studied global grokking—a synchronous memorization of most training data and a delayed contemporary generalization leap on most test samples—cannot be found in LLM pretraining. That being said, the memorization of certain data is plausible to impose a long-term, delayed effect on the generalization to specific downstream tasks. To investigate such local dynamic patterns, we group data in each domain by their memorization steps $t_{i}^{*}$ and particularly focus on the early-memorized groups, as they leave more checkpoints after $t_{i}^{*}$ to study the subsequent dynamics of generalization. Accordingly, we group samples in benchmarks by the training steps since when their predictions become consistently correct. To enable an analysis of how memorization drives downstream generalization performance, we pair each training data group with its most relevant benchmark samples via bipartite graph matching (by Hungarian algorithm [11]), with the training-test data similarity defined in the embedding space of SentenceTransformer [22].
Figure 2: Pretraining objective (blue) and benchmark performance (red) across four paired training-test data groups from the math domain. We observe local grokking on each pair: a delayed generalization leap on benchmark tasks after the pretraining objective plateaued. From left to right, more difficult data memorized (converged) later associate with a longer delay towards grokking.
@@ 4.1 Pathway Distance and Knowledge Sharing between Training Samples
We test this by examining how the similarity of expert pathway evolves during pretraining. Each input’s pathway is the ordered sequence of selected experts through the $L$ MoE layers. Specifically, for input $x_{i}$, we define its pathway $s_{i}$ as: $s_{i}=\mathrm{concat}(e_{1}^{(i)},\,e_{2}^{(i)},\,\dots,\,e_{L}^{(i)})$, where $e_{\ell}^{(i)}$ denotes the ordered list of expert indices at layer $\ell$, obtained by ranking experts according to the mean routing weights across all tokens for input $x_{i}$. To determine which experts are included, we first rank experts based on their routing weight and then select top-ranked experts until their cumulative routing weight exceeds a predefined threshold. The weight threshold selects the experts that actually drive each input’s computation, rather than picking a fixed number of experts. The selected experts are then concatenated as a comma-separated string (e.g., ‘3,1,5’) which are later joined across layers with hyphens (e.g., ‘3,1,5-…-9,1’).
@@
Layerwise pattern. Figure 5 reveals that the memorization-to-generalization transition is highly depth-dependent. Shallow layers show the steepest decline in $D_{\mathrm{path}}$ after memorization, indicating rapid convergence to shared pathways. Deeper layers exhibit smaller reductions, and the final layer even shows an increase in $D_{\mathrm{path}}$ post-memorization. These patterns suggest a depth-specific reorganization: early layers consolidate universal representations, while later layers retain or enhance routing flexibility, balancing shared computation with task-specific specialization.
@@ 4.2 Pathway Consistency and Complexity for Single Samples
where $g_{\ell}^{(i,k)}$ is the average routing weight assigned to expert $k$ across all tokens of input $x_{i}$. Although experts are independently parameterized, the strong correlations between adjacent layers in deep networks [6] and MoEs [12] suggest that $\{\mathbf{v}_{\ell}^{k}\}$ can be comparable across neighboring layers, supporting structural analysis of routing behavior.
@@ 4.3 Pathway Complexity Metrics as Monitors of Generalization
Given the strong link between routing dynamics and generalization, we ask: To what extent do our pathway metrics predict test performance across domains? To answer this, we compute Pearson and Spearman correlations between downstream performance and two metrics across pretraining checkpoints. Our correlation analysis focuses on checkpoints after domain-level accuracy begins its sustained rise (shown in Figure 1), thereby isolating the mechanisms that drive generalization beyond the random-accuracy phase.

### v4

@@ Abstract
This paper presents the first study of grokking in practical LLM pretraining. Specifically, we investigate when an LLM memorizes the training data, when its generalization on downstream tasks starts to improve, and what happens if there is a lag between the two. Unlike existing works studying when a small model generalizes to limited and specified tasks during thousands epochs’ training on algorithmic data, we focus on a practical setting for LLMs, i.e., near single-pass pretraining of next-token prediction on a cross-domain, large-scale corpus, and generalization on diverse benchmark tasks covering math/commonsense reasoning, code generation, and domain-specific retrieval. Our study, for the first time, verifies that grokking still emerges in pretraining mixture-of-experts (MoE) LLMs, though different local data groups may enter their grokking stages asynchronously due to the heterogeneity of their distributions and attributions to others. To find a mechanistic interpretation of this local grokking, we investigate the dynamics of training data’s pathways (i.e., expert choices across layers in MoE). Our primary discovery is that the pathways evolve from random, non-smooth across layers, instance-specific to more structured and transferable across samples, despite the converged pretraining loss. This depicts a transition from memorization to generalization. Two novel metrics are developed to quantify these patterns: one computes the pathway similarity between samples, while the other measures the consistency of aggregated experts between subsequent layers for each sample. These training data based metrics induce near zero cost but can faithfully track and monitor the generalization of LLMs on downstream tasks, reducing reliance on costly instruction tuning and benchmark evaluations. We also ground our findings in a theoretical analysis of one-layer MoE, showing that more structured pathways improve the generalization bound.
@@ 1 Introduction
  • Grokking still occurs during the near single-pass pretraining of practical-scale LLMs, but it is local and asynchronous for different data groups, unlike global grokking for all data in previous works.
@@ 3 Grokking (Delayed Generalization) in LLM Pretraining
  • Near single-pass pretraining: Most LLMs’ pretraining only takes approximately one epoch, so the model has been trained only once on each sample, and its training loss may even converge before visiting later data. However, such “memorization” of unseen data is not caused by overfitting or repeated replay. Instead, it reflects the interpolation or composition of learned data, which might be counted as “generalization” in previous works but differs from the emergent generalization of LLMs on new tasks.
@@ 3.1 Study Training Dynamics of LLM Pretraining
Since $\ell_{i}(\theta)$ directly measures how the model memorizes each token in $x_{i}$, we will track it in the pretraining process and identify the sample with consistently small and converged loss across checkpoints as memorized. Specifically, for a sample $x_{i}$ with loss $\ell_{i}(\theta_{t})$ at pretraining step $t$, we identify it as memorized since step $t_{i}^{*}$ if $\ell_{i}(\theta_{t})\leq\varepsilon$ and $|\ell_{i}(\theta_{t})-\ell_{i}(\theta_{T})|\leq\delta$ for all $t\geq t_{i}^{*}$, where $\varepsilon$ is a small threshold, $\delta$ enforces stability over time, and $T$ indexes the final pretraining checkpoint. In Appendix E.3, We verify that varying $\varepsilon$ and $\delta$ does not change the post-memorization behavior analyzed in Section 4.
@@ 3.2 Asynchronous Memorization, Delayed Generalization, & Local Grokking
These observations reveal the complex training dynamics of LLMs caused by training data heterogeneity and their uneven attributions to different benchmark tasks. It implies that the widely studied global grokking—a synchronous memorization of most training data and a delayed contemporary generalization leap on most test samples—cannot be found in LLM pretraining. That being said, the memorization of certain data is plausible to impose a long-term, delayed effect on the generalization to specific downstream tasks. To investigate such local dynamic patterns, we group data in each domain by their memorization steps $t_{i}^{*}$ and particularly focus on the early-memorized groups, as they leave more checkpoints after $t_{i}^{*}$ to study the subsequent dynamics of generalization. Accordingly, we group samples in benchmarks by the training steps since when their predictions become consistently correct. To enable an analysis of how memorization drives downstream generalization performance, we pair each training data group with its most relevant benchmark samples via bipartite graph matching (by Hungarian algorithm [13]), with the training-test data similarity defined in the embedding space of SentenceTransformer [27]. Additional implementation details of the pairing procedure are provided in Appendix B.1.
Figure 2: Pretraining objective (blue) and benchmark performance (red) across four paired training-test data groups from the math domain. Pretrain@$p$ → Test@$q$ denotes a paired training–test group, where Pretrain@$p$ is the training subset whose samples reach stable memorization at step $p$, and Test@$q$ is the benchmark subset whose accuracy stabilizes at step $q$. These pairs are formed using embedding-based Hungarian matching, as detailed in Appendix B.1. We observe local grokking on each pair: a delayed generalization leap on benchmark tasks after the pretraining objective plateaued. From left to right, more difficult data memorized (converged) later associate with a longer delay towards grokking.
@@ 4.1 Pathway Distance and Knowledge Sharing between Training Samples
We test this by examining how the similarity of expert pathway evolves during pretraining. Each input’s pathway is the ordered sequence of selected experts through the $L$ MoE layers. Specifically, for input $x_{i}$, we define its pathway $s_{i}$ as: $s_{i}=\mathrm{concat}(e_{1}^{(i)},\,e_{2}^{(i)},\,\dots,\,e_{L}^{(i)})$, where $e_{\ell}^{(i)}$ denotes the ordered list of expert indices at layer $\ell$, obtained by ranking experts according to the mean routing weights across all tokens for input $x_{i}$. To determine which experts are included, we first rank experts based on their routing weight and then select top-ranked experts until their cumulative routing weight exceeds a predefined threshold. The weight threshold selects the experts that actually drive each input’s computation, rather than picking a fixed number of experts. In our main analysis, we use a cumulative routing-weight threshold of 0.7, and Appendix E.1 shows that the observed pathway trends remain consistent when varying this threshold. The selected experts are then concatenated as a comma-separated string (e.g., ‘3,1,5’) which are later joined across layers with hyphens (e.g., ‘3,1,5-…-9,1’).
@@
Layerwise pattern. While prior work [17] reports static depth-dependent specialization in trained MoEs, our results show that the memorization-to-generalization transition is also dynamically depth-dependent during pretraining (Figure 5). Shallow layers show the steepest decline in $D_{\mathrm{path}}$ after memorization, indicating rapid convergence to shared pathways. Deeper layers exhibit smaller reductions, and the final layer even shows an increase in $D_{\mathrm{path}}$ post-memorization. These patterns suggest a depth-specific reorganization: early layers consolidate universal representations, while later layers retain or enhance routing flexibility, balancing shared computation with task-specific specialization.
@@ 4.2 Pathway Consistency and Complexity for Single Samples
where $g_{\ell}^{(i,k)}$ is the routing weight assigned to expert $k$. Although experts are independently parameterized, the strong correlations between adjacent layers in deep networks [8] and MoEs [14] suggest that $\{\mathbf{v}_{\ell}^{k}\}$ can be comparable across neighboring layers, supporting structural analysis of routing behavior.
@@ 4.3 Pathway Complexity Metrics as Monitors of Generalization
Given the strong link between routing dynamics and generalization, we ask: To what extent do our pathway metrics predict test performance across domains? To answer this, we compute Pearson and Spearman correlations between downstream performance and two metrics across pretraining checkpoints. As described in Section 3.1, we apply a lightweight LoRA-based instruction tuning (rank = 32; see Appendix B) solely to equip pretrained checkpoints with basic instruction-following ability for benchmark evaluation. The robustness of this setup is further supported by the LoRA rank ablation in Appendix E.2. Our correlation analysis focuses on checkpoints after domain-level accuracy begins its sustained rise (shown in Figure 1), thereby isolating the mechanisms that drive generalization beyond the random-accuracy phase.
@@ Appendix B Experiment Details
B.1 Construction of Training–Test Paired Groups.
Here we provide additional details on how we construct the training–test paired groups used in Figure 2 to analyze the temporal relation between memorization and generalization. The overall procedure is introduced in Section 3.2; here, we summarize the specific grouping and matching setup for clarity and reproducibility.
Training groups.
[…]
Test groups.
[…]
Embedding-based pairing.
To align semantically related training and test groups, we compute sentence embeddings for all samples using the SentenceTransformer model “all-MiniLM-L6-v2” and measure pairwise cosine similarity between groups within each domain. The average similarity between every pair of groups forms a similarity matrix, and a one-to-one mapping between training and test groups is obtained using the Hungarian algorithm implemented in SciPy, which maximizes overall semantic similarity. The notation Pretrain@$p$ $\rightarrow$ Test@$q$ thus indicates that the training group memorized at step $p$ is semantically aligned with a test group whose performance stabilizes at step $q$.
@@ Results.
Appendix E Analysis
E.1 Effect of Routing-Weight Threshold on Pathway Dynamics
To examine the robustness of our findings with respect to the routing-weight threshold used for pathway construction, we repeat the analysis in Fig. 4 under different cumulative routing-weight thresholds ranging from 0.5 to 0.8 (interval 0.1). This threshold determines how many experts are considered active in each layer when forming the pathway sequence: a higher threshold includes more experts, while a lower threshold selects only the most dominant ones.
Figure 9: Effect of routing-weight threshold on pathway edit distance on domain-specific QA tasks. Each panel shows the average pathway edit distance for this domain under a different cumulative routing-weight threshold (0.5–0.8, interval 0.1). Higher thresholds include more active experts, leading to longer absolute distances, but all settings exhibit the same qualitative pattern: pathway edit distance decreases consistently after memorization, confirming that the observed dynamics are robust to the threshold choice.
Figure 9 presents the results for the domain-specific QA tasks. Although the absolute values of edit distance increase with larger thresholds—reflecting the longer sequences formed when more experts are included—the qualitative trends remain consistent across all settings. For every threshold, the pathway edit distance decreases after memorization (marked by vertical dashed lines), and the decrease becomes more pronounced at later checkpoints.
These results demonstrate that our conclusions are robust to the choice of routing-weight threshold. The observed reduction in pathway complexity after memorization, corresponding to the emergence of shared pathways, persists across all threshold settings.
E.2 Effect of LoRA Rank on Pathway–Generalization Correlation
To examine whether the correlation between pathway metrics and generalization depends on the configuration of instruction tuning, we conduct an ablation on the LoRA rank used during lightweight finetuning. As described in Appendix B, the default setup adopts a rank of 32 for all domains. Here, we additionally train a LoRA adapter with rank 64 on the commonsense domain, following the same data, training steps, and hyperparameters as in the main experiments.
Table 4 reports the correlation between pathway complexity metrics and downstream test accuracy under the two LoRA ranks. The results show that increasing the rank from 32 to 64 yields nearly identical correlations for both Pathway Distance and Pathway Consistency, while training and validation losses remain weakly correlated. This consistency indicates that the observed relationships between pathway metrics and generalization are robust to LoRA configuration and primarily reflect intrinsic pretraining-phase dynamics rather than artifacts of instruction tuning.
Table 4: Effect of LoRA Rank on the Correlation Between Pathway Metrics and Generalization (Commonsense reasoning domain). The results show that increasing the LoRA rank from 32 to 64 yields nearly identical correlations, confirming that the observed trends are robust to LoRA configuration. $p$-value: $p<0.1$, $p<0.05$, $p<0.01$, $p>0.1$, $p>0.5$, $p>0.8$. Preferred correlation: positive, negative.
Metric LoRA Rank = 32 LoRA Rank = 64
Pearson Spearman Pearson Spearman
[…]
E.3 Effect of Memorization Thresholds $\varepsilon$ and $\delta$
We conduct an ablation study on the two hyperparameters that define the memorization criterion in Section 3.1: the loss threshold $\varepsilon$ and the stability threshold $\delta$. Specifically, for a sample $x_{i}$ with loss $\ell_{i}(\theta_{t})$ at pretraining step $t$, it is considered memorized from step $t_{i}^{*}$ if $\ell_{i}(\theta_{t})\leq\varepsilon$ and $|\ell_{i}(\theta_{t})-\ell_{i}(\theta_{T})|\leq\delta$ for all $t\geq t_{i}^{*}$. We evaluate how sensitive the pathway trends are to these two parameters using the domain-specific QA task.
Varying $\varepsilon$.
Figure 10 shows results when fixing $\delta=0.05$ and varying $\varepsilon$ from 1.5 to 2.0. Across all settings, the pathway edit distance after memorization consistently decreases with training, and the overall trend remains stable—indicating that $\varepsilon$ only shifts the point where samples are considered memorized but does not affect the post-memorization dynamics.
Varying $\delta$.
Figure 11 reports the results when fixing $\varepsilon=2.0$ and varying $\delta$ between 0.03 and 0.05. Similarly, the decreasing trend of edit distance after memorization persists, and the fitted trajectories show negligible variation across $\delta$ values.
This confirms that the proposed pathway-based observations are robust to the choice of $\varepsilon$ and $\delta$.
Figure 10: Effect of the loss threshold $\varepsilon$ on pathway dynamics. The stability threshold is fixed at $\delta=0.05$. All curves show a consistent post-memorization decrease in edit distance, indicating that the results are robust to $\varepsilon$.
Figure 11: Effect of the stability threshold $\delta$ on pathway dynamics. The loss threshold is fixed at $\varepsilon=2.0$. The same decreasing trend holds across different $\delta$ values, confirming robustness to this parameter.
E.4 Cross-Domain Analysis of Pathway Dynamics
To verify that the observed changes in pathway edit distance are due to the emergence of shared pathways within specific domains rather than uniform effects across all data, such as load balancing or regularization, we conduct an additional analysis focusing on the cross-domain samples. Using the same cumulative routing-weight threshold (0.7) as in the main experiments, we compute average pathway edit distances for samples drawn from two different domains across checkpoints—specifically, math $\times$ code—under identical configurations.
Figure 12: Cross-domain pathway edit distance (math $\times$ code). The figure shows the global trend of average pathway edit distance during pretraining when computed across two different domains. The cross-domain distances fluctuate irregularly, exhibiting both increases and decreases but no systematic downward trend. This contrasts with the localized decreases observed within domains (Fig. 4), indicating that pathway reorganization occurs locally among semantically related data rather than globally across domains.
As shown in Fig. 12, the cross-domain edit distance fluctuates throughout pretraining, showing irregular increases and decreases without any consistent downward trend. In contrast, Fig. 4 demonstrates that within-domain pathway distances decrease only for groups of samples that become memorized at different points in training, with each domain exhibiting its own localized transition rather than a global drop. These observations confirm that the structural changes we observe are domain- and group-specific, reflecting the emergence of shared pathways among semantically related samples rather than a global effect of routing behavior.
E.5 Relationship Between Routing Entropy and Edit Distance
We investigate whether changes in routing diversity could trivially explain the observed trends in pathway edit distance. Specifically, we test whether routing entropy—high when expert usage is balanced and low when a few experts dominate—can drive edit distance dynamics.
For each sample, routing entropy is computed per layer from the full routing-weight vector (no thresholding or selection) and averaged across layers. For each sample pair, we take the mean of the two samples’ entropies and compute their pathway edit distance using a cumulative routing-weight threshold of 0.8, which provides sufficient expert coverage for analysis.
We sample 141,530 pairs across checkpoints spanning both the math and code domains, including within-domain and cross-domain pairs. Figure 13 shows a weak correlation between average routing entropy and edit distance ($r=0.24$, explaining about $6\%$ of the variance) with no clear monotonic or directional trend. This indicates that neither more balanced routing (higher entropy) nor more concentrated routing (lower entropy) systematically reduces pathway distance. Therefore, the decrease in within-domain pathway distance observed after memorization is unlikely to result from overall changes in routing balance; instead, it reflects domain- and group-specific reorganization of computation.
Figure 13: Edit distance vs. routing entropy. Each hexagon aggregates sample pairs (N=141,530N=141{,}530) drawn across checkpoints and across the math and code domains. Routing entropy is computed from the full routing-weight vectors (no thresholding); pathway edit distance uses pathways built with a cumulative cutoff of $0.8$. The weak correlation ($r=0.24$) indicates that routing balance does not systematically determine pathway similarity.
E.6 Consistent Findings on another MoE (nanoMoE)
To further validate the generality of our findings beyond OLMoE, we conduct additional experiments on a 55M-parameter MoE model, nanoMoE [^1] https://github.com/wolfecameron/nanoMoE. NanoMoE provides a small, fully trainable MoE architecture for which we can obtain complete pretraining checkpoints, making it suitable for analysis of routing dynamics.
Training setup. We train nanoMoE on $\sim$25B tokens of the OpenWebText dataset [7] for 50k iterations on two NVIDIA H100 GPUs and uniformly sample ten checkpoints for analysis. OpenWebText is a mixed-domain web corpus with a heavy emphasis on commonsense and general knowledge content. To approximate a domain setting, we construct a commonsense subset by filtering 10,000 randomly sampled training samples using Qwen3-8B [33] to select those semantically related to commonsense reasoning, yielding 7,857 samples in total. Following the same methodology as our main experiments, we identify memorized samples based on token-level loss stability and compute pathway dynamics for this subset across checkpoints.
Pathway-metric dynamics. Consistent with our large-scale results, we observe that after memorization, pathway edit distance decreases while pathway consistency increases as shown in Figure 14 and  15. These trends mirror the post-memorization reorganization we identify in OLMoE, demonstrating that even the smaller MoE model reorganizes routing toward more structured, shared computation after the loss has already converged.
Figure 14: Pathway edit distance and pretraining objective on nanoMoE during pretraining. The left panel reports the overall pathway edit-distance trajectory across checkpoints without grouping samples by their memorization steps. The right panel instead visualizes pathway dynamics for two memorized data groups (memorized at 30k and 45k steps). After their respective memorization points, both groups exhibit a consistent post-memorization decline in edit distance despite a plateaued pretraining objective, indicating that nanoMoE continues to reorganize routing pathways toward more structured and shared computation. This mirrors the memorization-to-generalization transition observed in large-scale OLMoE.
Figure 15: Evolution of pathway consistency in nanoMoE. The left panel reports the overall progression of pathway consistency across checkpoints. The right panel further analyzes two memorized data groups (memorized at 30k and 35k steps), showing that once memorization is achieved, each group undergoes continued refinement of its layer-to-layer routing transitions.
Correlation with downstream performance. We further evaluate all ten checkpoints in a zero-shot setting on standard commonsense reasoning benchmarks (ARC-Easy/Challenge [4], HellaSwag [34], and OpenBookQA [20]). Zero-shot evaluation is justified because models pretrained on OpenWebText, including GPT-2 [26], have been shown to achieve strong zero-shot performance on commonsense benchmarks. As in the large-scale setting, both pathway metrics exhibit strong correlations with benchmark accuracy ($|r|\approx 0.8$$1.0$), as reported in Table 5, substantially outperforming training loss or its moving average. This confirms that routing dynamics remain tightly coupled with downstream generalization.
The nanoMoE results reinforce our central claim: post-memorization routing reorganization is not an artifact of model scale, data size, or specific MoE implementation. Instead, it reflects an inherent property of MoE training dynamics. Combined with the large-scale OLMoE study, these findings establish pathway metrics as stable, architecture-intrinsic indicators of the memorization-to-generalization transition.
Table 5: Correlation Between Pathway Metrics and Downstream Accuracy on nanoMoE across four commonsense benchmarks (ARC-Easy, ARC-Challenge, HellaSwag, OpenBookQA). $p$-value: $p<0.1$, $p<0.05$, $p<0.01$, $p>0.1$, $p>0.5$, $p>0.8$. Preferred correlation: positive, negative.
Metric ARC-Easy ARC-Challenge HellaSwag OpenBookQA
[…]
LLM Usage Statement
In preparing this manuscript, we used LLM solely as a writing assistance tool. Specifically, the model was employed to aid in grammar correction, sentence refinement, and improving clarity of exposition. It did not contribute to research ideation, experimental design, data analysis, or interpretation of results.